OpenStudy (adamlee):

What is the acceleration of a ball traveling horizontally with an initial velocity of 20 meters/second and, 2.0 seconds later, a velocity of 30 meters/second?

OpenStudy (logo):

Vf = Vi + at Vf is final and Vi is initial velocity 30 = 20 + a(2) a = 5 m/s^2